# MeanSparse: Post-Training Robustness Enhancement Through Mean-Centered Feature Sparsification

### v2

@@ Abstract
We present a simple yet effective method to improve the robustness of both Convolutional and attention-based Neural Networks against adversarial examples by post-processing an adversarially trained model. Our technique, MeanSparse, cascades the activation functions of a trained model with novel operators that sparsify mean-centered feature vectors. This is equivalent to reducing feature variations around the mean, and we show that such reduced variations merely affect the model’s utility, yet they strongly attenuate the adversarial perturbations and decrease the attacker’s success rate. Our experiments show that, when applied to the top models in the RobustBench leaderboard, MeanSparse achieves a new robustness record of $75.28\%$ (from $73.71\%$), $44.78\%$ (from $42.67\%$) and $62.12\%$ (from $59.56\%$) on CIFAR-10, CIFAR-100 and ImageNet, respectively, in terms of AutoAttack accuracy. Code: https://github.com/SPIN-UMass/MeanSparse
1 Introduction
@@
MeanSparse establishes a new state-of-the-art (SOTA) in robustness. When applied to the leading robust models, it improves $\ell_{\infty}$ AutoAttack accuracy from $73.71\%$ to $75.28\%$ on CIFAR-10, from $42.67\%$ to $44.78\%$ on CIFAR-100, and from $59.56\%$ to $62.12\%$ on ImageNet, while maintaining nearly unchanged clean accuracy. Additionally, when applied to the top model for $\ell_{2}$ AutoAttack accuracy on CIFAR-10, MeanSparse boosts robust accuracy from $84.97\%$ to $87.28\%$. In all four cases, MeanSparse achieves the highest ranking according to RobustBench specifications [9].
Figure 1: Mean-based sparsification operator used in the MeanSparse technique for hypothetical channel ch. The first column represents the design procedure. First, the mean ($\mu_{\mathrm{ch}}$) and standard deviation ($\sigma_{\mathrm{ch}}$) are calculated over the training set (top figure). The mean-based sparsification operator is designed with hyper-parameter $\alpha$ which blocks the variations in the $\alpha\sigma_{\mathrm{ch}}$ vicinity of $\mu_{\mathrm{ch}}$ (bottom figure). The second column represents how mean-based sparsification affects the input features for one test sample (top figure) and generates output features (bottom figure). The effect of mean-based sparsification over the feature histogram is also demonstrated in the third column.
@@
2 Preliminaries
2.1 Notations
We use bold lowercase letters to show vectors while the $i$-th element of vector $\bm{x}$ is represented by $x_{i}$. Two vector norm metrics are used throughout the paper, including $\ell_{0}=\|\cdot\|_{0}$ (number of non-zero elements) and $\ell_{2}=\|\cdot\|_{2}$ (Euclidean norm). The $\ell_{0}$ norm can also effectively penalize the variables with small values but its nonsmoothness limits its application. The proximal operator finds the minimizer of function $f$ in the vicinity of input vector $\bm{v}$ and plays an essential role in our intuition to design the mean-based sparsification operator.
Definition 2.1 (Proximal operator [33]).
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a proper and lower semi-continuous function. The proximal operator (mapping) $\mathop{\mathrm{prox}}_{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of $f$ at $\bm{v}$ is defined as:
$\mathop{\mathrm{prox}}_{f}(\bm{v})=\operatorname*{arg\,min}_{\bm{x}}~{}f(\bm{x})+\frac{1}{2}\|\bm{x}-\bm{v}\|_{2}^{2}$ (1)
The definition of proximal operator can be extended to nonsmooth functions such as $\ell_{0}$ norm. Assume $f(\bm{a})\triangleq\lambda\|\bm{a}\|_{0}$. Then the proximal operator can be calculated as [13]:
$\mathop{\mathrm{prox}}_{f}(\bm{v})=\mathcal{H}_{2\lambda}(\bm{v}),~{}\mathcal{H}_{\alpha}(v)=\begin{cases}v&|v|>\sqrt{\alpha}\\ 0&|v|\leq\sqrt{\alpha}\end{cases}$ (2)
@@ 2.2 Related Work
Parametric activation functions increase the flexibility of activation functions while the parameters can be learned during the training phase. This flexibility, if designed intelligently, is also shown to be effective in improving robustness with AT. An example of Parametric activation functions is parametric shifted SiLU (PSSILU) [11].
Similar to Neural Architecture Search (NAS), Searching for Adversarial Robust Activation Functions (SARAF) is designed to search over candidate activation functions to maximize the model robustness. Using a meta-heuristic search method, SARAF can handle the intractable complexity of the original search problem [39, 38].
3 Methodology
@@
3.1 Intuition from Regularized Optimization Objective
@@
Figure 2: Mean-centered feature used in regularized optimization problem equation 3
@@
The small changes around the feature mean value at the output of different layers are potentially non-robust features. Thus we start by penalizing those changes using $\ell_{0}$ norm in an arbitrary layer $l$ for a training sample $\bm{x}$ (extension to the whole dataset and all layers is straightforward). The original training objective can be formulated as:
$\mathcal{P}_{0}:~{}\bm{\theta}_{0}^{\star}=$ $\operatorname*{arg\,min}_{\bm{\theta}}~{}L(\bm{\theta})+\gamma\|\overline{\bm{a}}(\bm{x};\bm{\theta})\|_{0},~{}~{}L(\bm{\theta})\triangleq\mathcal{L}(\bm{y}(\bm{x}),\bm{y}^{\star};\bm{\theta})$ (3)
@@
where $\bm{\theta}$ is the vector containing model parameters, $\gamma$ is a hyper-parameter, $\overline{\bm{a}}(\bm{x};\bm{\theta})$ is the centralized feature vector before the activation function in the regularized layer, $\bm{y}(\cdot)$ is the model output and $\bm{y}^{\star}$ is the target label corresponding to $\bm{x}$. Throughout the paper, we assume $\mu_{a}$ to be a fixed vector that is updated using an exponential filter during training, similar to the mean and variance vectors calculated in a batch normalization layer. Figure 2 represents the way mean-centered features for regularization are generated. One option to solve the optimization problem in equation 3 is via the penalty method.
@@
To solve problem equation 3, we introduce an approximate version of this problem based on the penalty method (for simplicity we use $\overline{\bm{a}}$ instead of $\overline{\bm{a}}(\bm{x};\bm{\theta})$):
@@
$\mathcal{P}_{\lambda}:~{}\bm{\theta}^{\star}_{0,\lambda},\bm{w}^{\star}=$ $\operatorname*{arg\,min}_{\bm{\theta},\bm{w}}~{}L(\bm{\theta})+\gamma\|\bm{w}\|_{0}+\frac{1}{2\lambda}\|\bm{w}-\overline{\bm{a}}\|_{2}^{2}$ (4)
where $\lambda$ is the penalty parameter, $\bm{w}$ is the penalty variable, and the optimization is with respect to both variables $\bm{\theta}$ and $\bm{w}$. Based on the penalty method (PM), we can conclude:
$\lim_{\lambda\rightarrow 0}\bm{\theta}_{0,\lambda}^{\star}=\bm{\theta}^{\star}_{0}$ (5)
@@
So to approximate $\bm{\theta}_{0}^{\star}$, one can solve problem $\mathcal{P}_{\lambda}$ and decrease $\lambda$ during the training. One approach to solve problem $\mathcal{P}_{\lambda}$ is to use block coordinate descent where the $\bm{\theta}$ and $\bm{w}$ are updated sequentially [31]. So we have two steps in each training iteration $k$.
Step1$\rightarrow$ Calculatign $\bm{w}_{k}$: To calculate $\bm{w}_{k}$, we should solve:
$\bm{w}_{k}$ $=\operatorname*{arg\,min}_{\bm{w}}~{}\gamma\lambda\|\bm{w}\|_{0}+\frac{1}{2}\|\bm{w}-\overline{\bm{a}}_{k-1}\|_{2}^{2}$ (6)
@@
The problem in equation 6 matches the definition of the proximal operator in equation 1 and can be solved exactly by the proximal operator for the function $f(\bm{w})=\gamma\lambda\|\bm{w}\|_{0}$. Thus we have:
@@
$\bm{w}_{k}=\mathop{\mathrm{prox}}_{f}(\bm{z}_{k-1})=\mathcal{H}_{2\lambda\gamma}(\overline{\bm{a}}_{k-1})$ (7)
@@
Step 2 $\rightarrow$ Calculating $\bm{\theta}_{k}$: To calculate $\bm{\theta}_{k}$, we have to solve the following problem:
$\bm{\theta}_{k}=\operatorname*{arg\,min}_{\bm{\theta}}~{}L(\bm{\theta})+\frac{1}{2\lambda}\|\bm{w}_{k}-\overline{\bm{a}}\|_{2}^{2}$ (8)
@@
while $\bm{w}_{k}$ is calculated using hard-thresholding operator as in equation 7. One step of (stochastic) gradient descent can be used to update $\bm{\theta}$ based on equation 8.
3.2 Mean-centered Feature Sparsification
@@
Update rule equation 8 shows $\overline{\bm{a}}$ approaches $\mathcal{H}_{2\lambda\gamma}(\overline{\bm{a}}_{k-1})$ by the second term while the weight for this term is increasing (equivalently $\lambda$ is decreasing) during the training. Based on this intuition, we propose to use the $\mathcal{H}_{2\lambda\gamma}(\overline{\bm{a}}_{k-1})$ in the forward propagation of the model. If we add the element-wise mean subtraction and addition in 2 into the hard-thresholding operator, we achieve the curve in the bottom of the first column in 1 assuming $\alpha\sigma_{\mathrm{ch}}=(2\lambda\gamma)^{2}$ (for simplicity of notation, we use $Th$ instead). This curve represents a sparsification around the feature mean value $\mu_{a}$. In other words, we have the following element-wise operation over the input $\bm{a}^{(\text{in})}$:
$a^{(\text{out})}=\begin{cases}\mu_{a}&if~{}|a^{(\text{out})}-\mu_{a}|\leq~{}Th\\ a^{(\text{in})}&if~{}|a^{(\text{out})}-\mu_{a}|>~{}Th\\ \end{cases}$
@@ 3.3 MeanSparse Technique
Sparsification in Post-processing
@@
In this paper, we use the second method. Thus, we freeze the model and apply sparsification over mean-centered features before the activation functions in the trained model. Then calculate the mean value over the training set and finally find the value of $Th$ that results in the best robustness performance while maintaining the model’s clean accuracy.
Adaptive Sparsification Using Feature Standard Deviation
@@
Per-channel Sparsification
Throughout the neural network, the input to the activation functions can be considered as 4-dimensional, represented as $N\times C\times H\times W$ ($N$, $C$, $H$ and $W$ represents the batch size, channels, height and width, respectively). To better capture the statistics in the representation, we use per-channel mean $\mu_{\mathrm{ch}}$ and variance $\sigma_{\mathrm{ch}}$ in sparsification operator. As a result, the mean and variance vectors are $C$-dimensional vectors. The sparsification operates on each feature channel separately, using the mean and variance of the corresponding channel.
Complete Pipeline
The complete pipeline to apply MeanSparse technique is formed by equipping the mean-centered sparsification operator with the three aforementioned modifications and adding it before all activation functions in the model with a shared $\alpha$ parameter. Figure 1 illustrates an example of integrating the MeanSparse technique into an adversarially trained model. In the first column, the upper figure represents the histogram of one channel features in the first layer of RaWideResNet trained adversarially over the CIFAR-10 dataset [34] along with the mean $\mu_{\mathrm{ch}}$ and standard deviation $\sigma_{\mathrm{ch}}$ over the complete training set. Based on this histogram, we design the bottom sparsification operator where the variations in $\alpha\sigma_{\mathrm{ch}}$-vicinity of the mean are blocked. The second and third column represents the input and output of the sparsification operator for one test image of the CIFAR-10 dataset $\bm{x}_{t}$ in sample space and histogram, respectively. The top figures are input and the bottom ones are output. The feature histogram for $\bm{x}_{t}$ is bimodal. One of the modes matched the $\mu_{\mathrm{ch}}$ which is the feature’s peak over the training set. Based on our intuition, the variations in this region are uninformative and can be blocked as we see in the bottom figures of the second and third columns.

### v3

@@ Abstract
We present a simple yet effective method to improve the robustness of both Convolutional and attention-based Neural Networks against adversarial examples by post-processing an adversarially trained model. Our technique, MeanSparse, cascades the activation functions of a trained model with novel operators that sparsify mean-centered feature vectors. This is equivalent to reducing feature variations around the mean, and we show that such reduced variations merely affect the model’s utility, yet they strongly attenuate the adversarial perturbations and decrease the attacker’s success rate. Our experiments show that, when applied to the top models in the RobustBench leaderboard, MeanSparse achieves a new AutoAttack robust accuracy record of $75.28\%$ (from $73.71\%$), $44.78\%$ (from $42.67\%$) and $62.12\%$ (from $59.56\%$) on CIFAR-10, CIFAR-100 and ImageNet, respectively, in terms of AutoAttack accuracy. Code: https://github.com/SPIN-UMass/MeanSparse
I Introduction
@@
The MeanSparse operator selectively suppresses variations around the mean of feature representations, effectively filtering out non-robust features. For a given feature channel, we compute the mean ($\mu$) and standard deviation ($\sigma$) over the training set. Using a tunable threshold ($Th=\alpha\sigma$), we block feature values that lie within $\mu\pm Th$, replacing them with the mean value ($\mu$). This operation limits minor perturbations that adversarial attacks often exploit, while preserving the informative structure of features outside this range. For instance, consider a hypothetical feature channel with a mean ($\mu$) of 0.5 and standard deviation ($\sigma$) of 0.2. Setting $\alpha=1$, we block values between 0.3 and 0.7, replacing them with 0.5. This simple mechanism attenuates insignificant variations, as demonstrated in Figure 1 of the paper, where we visualize how the input histogram is transformed. The blocked region corresponds to low-information variations, enhancing robustness by reducing the attacker’s exploitable capacity.
Figure 1: Mean-based sparsification operator used in the MeanSparse technique for hypothetical channel ch. The first column represents the design procedue. First, the mean ($\mu_{\mathrm{ch}}$) and standard deviation ($\sigma_{\mathrm{ch}}$) are calculated over the training set (top figure). The mean-based sparsification operator is designed with hyper-parameter $\alpha$ which blocks the variations in the $\alpha\sigma_{\mathrm{ch}}$ vicinity of $\mu_{\mathrm{ch}}$ (bottom figure). The second column represents how mean-based sparsification affects the input features for one test sample (top figure) and generates output features (bottom figure). The effect of mean-based sparsification over the feature histogram is also demonstrated in the third column.
@@
MeanSparse establishes a new state-of-the-art (SOTA) in robustness in terms of AutoAttack accuracy. When applied to the leading robust models, it improves $\ell_{\infty}$ AutoAttack accuracy from $73.71\%$ to $75.28\%$ on CIFAR-10, from $42.67\%$ to $44.78\%$ on CIFAR-100, and from $59.56\%$ to $62.12\%$ on ImageNet, while maintaining nearly unchanged clean accuracy. Additionally, when applied to the top model for $\ell_{2}$ AutoAttack accuracy on CIFAR-10, MeanSparse boosts robust accuracy from $84.97\%$ to $87.28\%$. In all four cases, MeanSparse achieves the highest ranking in terms of AutoAttack accuracy [16].
@@
II Preliminaries
II-A Notations
We use bold lowercase letters to show vectors while the $i$-th element of vector $\boldsymbol{x}$ is represented by $x_{i}$. Two vector norm metrics are used throughout the paper, including $\ell_{0}=\|\cdot\|_{0}$ (number of non-zero elements) and $\ell_{2}=\|\cdot\|_{2}$ (Euclidean norm). The $\ell_{0}$ norm can also effectively penalize the variables with small values but its nonsmoothness limits its application. The proximal operator finds the minimizer of function $f$ in the vicinity of input vector $\boldsymbol{v}$ and plays an essential role in our intuition to design the mean-based sparsification operator.
Definition 1 (Proximal operator [17])
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a proper and lower semi-continuous function. The proximal operator (mapping) $\mathop{\mathrm{prox}}_{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of $f$ at $\boldsymbol{v}$ is defined as:
$\mathop{\mathrm{prox}}_{f}(\boldsymbol{v})=\operatornamewithlimits{arg\,min}_{\boldsymbol{x}}~f(\boldsymbol{x})+\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{v}\|_{2}^{2}$ (1)
The definition of proximal operator can be extended to nonsmooth functions such as $\ell_{0}$ norm. Assume $f(\boldsymbol{a})\triangleq\lambda\|\boldsymbol{a}\|_{0}$. Then the proximal operator can be calculated as [18]:
$\mathop{\mathrm{prox}}_{f}(\boldsymbol{v})=\mathcal{H}_{2\lambda}(\boldsymbol{v}),~\mathcal{H}_{\alpha}(v)=\begin{cases}v&|v|>\sqrt{\alpha}\\ 0&|v|\leq\sqrt{\alpha}\end{cases}$ (2)
@@ II-B Related Work
Parametric activation functions increase the flexibility of activation functions while the parameters can be learned during the training phase. This flexibility, if designed intelligently, is also shown to be effective in improving robustness with AT. An example of Parametric activation functions is parametric shifted SiLU (PSSILU) [9]. Similar to Neural Architecture Search (NAS), Searching for Adversarial Robust Activation Functions (SARAF) is designed to search over candidate activation functions to maximize the model robustness. Using a meta-heuristic search method, SARAF can handle the intractable complexity of the original search problem [44, 10].
II-C Threat Model
In this experiment, we focused on the effectiveness of MeanSparse against two well-studied threat models, $\ell_{\infty}$ and $\ell_{2}$. Until now, our focus was on the $\ell_{\infty}$ threat model, which generally results in lower robust accuracy [16]. On the other hand, $\ell_{2}$ threat models can focus on a specific part of the image, and this localization of adversarial perturbation may challenge the effectiveness of the MeanSparse technique. In this experiment, we compare the robustness against both $\ell_{\infty}$ and $\ell_{2}$ threat models. The model is ResNet-18 with the GELU activation function as in our Activation Function experiment in table II. Table III illustrates the clean and APGD accuracy for $\ell_{\infty}$ and $\ell_{2}$ threats. As the accuracy for the $\ell_{\infty}$ threat model increases with increasing $\alpha$, the accuracy for $\ell_{2}$ also increases, which depicts the effectiveness and generalization of MeanSparse across well-studied threat models. Additionally, the improvement compared to the base model at $\alpha=0.2$ is larger for the $\ell_{2}$ threat than for the $\ell_{\infty}$ threat.
III Methodology
@@
III-A Intuition from Regularized Optimization Objective
@@
Figure 2: Mean-centered feature used in regularized optimization problem of (3)
@@
The small changes around the feature mean value at the output of different layers are potentially non-robust features. Thus we start by penalizing those changes using $\ell_{0}$ norm in an arbitrary layer $l$ for a training sample $\boldsymbol{x}$ (extension to the whole dataset and all layers is straightforward). The original training objective can be formulated as:
$\mathcal{P}_{0}:~\boldsymbol{\theta}_{0}^{\star}$ $=\operatornamewithlimits{arg\,min}_{\boldsymbol{\theta}}~L(\boldsymbol{\theta})+\gamma\|\overline{\boldsymbol{a}}(\boldsymbol{x};\boldsymbol{\theta})\|_{0},$
$L(\boldsymbol{\theta})$ $\triangleq\mathcal{L}(\boldsymbol{y}(\boldsymbol{x}),\boldsymbol{y}^{\star};\boldsymbol{\theta})$ (3)
@@
where $\boldsymbol{\theta}$ is the model parameters vector, $\gamma$ is a hyper-parameter, $\overline{\boldsymbol{a}}(\boldsymbol{x};\boldsymbol{\theta})$ is the centralized feature vector before the activation function in the regularized layer, $\boldsymbol{y}(\cdot)$ is the model output and $\boldsymbol{y}^{\star}$ is the target label corresponding to $\boldsymbol{x}$. Throughout the paper, we assume $\mu_{a}$ to be a fixed vector that is updated using an exponential filter during training, similar to the mean and variance vectors calculated in a batch normalization layer. Figure 2 represents the way mean-centered features for regularization are generated. The optimization problem in (3) can be solved via penalty method.
@@
To solve problem (3), we introduce an approximate version of this problem based on the penalty method (for simplicity we use $\overline{\boldsymbol{a}}$ instead of $\overline{\boldsymbol{a}}(\boldsymbol{x};\boldsymbol{\theta})$):
@@
$\mathcal{P}_{\lambda}:~\boldsymbol{\theta}^{\star}_{0,\lambda},\boldsymbol{w}^{\star}=$ $\operatornamewithlimits{arg\,min}_{\boldsymbol{\theta},\boldsymbol{w}}~L(\boldsymbol{\theta})+\gamma\|\boldsymbol{w}\|_{0}+\frac{1}{2\lambda}\|\boldsymbol{w}-\overline{\boldsymbol{a}}\|_{2}^{2}$ (4)
@@
where $\lambda$ is the penalty parameter, $\boldsymbol{w}$ is the penalty variable, and the optimization is with respect to both variables $\boldsymbol{\theta}$ and $\boldsymbol{w}$. Based on the penalty method (PM), we can conclude: $\lim_{\lambda\rightarrow 0}\boldsymbol{\theta}_{0,\lambda}^{\star}=\boldsymbol{\theta}^{\star}_{0}$. So to approximate $\boldsymbol{\theta}_{0}^{\star}$, one can solve problem $\mathcal{P}_{\lambda}$ and decrease $\lambda$ during the training. One approach to solve problem $\mathcal{P}_{\lambda}$ is to use block coordinate descent where the $\boldsymbol{\theta}$ and $\boldsymbol{w}$ are updated sequentially [46]. So we have two steps in each training iteration $k$.
[…]
$\boldsymbol{w}_{k}$ $=\operatornamewithlimits{arg\,min}_{\boldsymbol{w}}~\gamma\lambda\|\boldsymbol{w}\|_{0}+\frac{1}{2}\|\boldsymbol{w}-\overline{\boldsymbol{a}}_{k-1}\|_{2}^{2}$ (5)
@@
The problem in (5) matches the definition of the proximal operator in (1) and can be solved exactly by the proximal operator for the function $f(\boldsymbol{w})=\gamma\lambda\|\boldsymbol{w}\|_{0}$. Thus we have:
@@
$\boldsymbol{w}_{k}=\mathop{\mathrm{prox}}_{f}(\bar{\boldsymbol{a}}_{k-1})=\mathcal{H}_{2\lambda\gamma}(\overline{\boldsymbol{a}}_{k-1})$ (6)
@@
Step 2 $\rightarrow$ Calculating $\boldsymbol{\theta}_{k}$: To calculate $\boldsymbol{\theta}_{k}$, we have to solve the following problem:
$\boldsymbol{\theta}_{k}=\operatornamewithlimits{arg\,min}_{\boldsymbol{\theta}}~L(\boldsymbol{\theta})+\frac{1}{2\lambda}\|\boldsymbol{w}_{k}-\overline{\boldsymbol{a}}\|_{2}^{2}$ (7)
@@
while $\boldsymbol{w}_{k}$ is calculated using hard-thresholding operator as in (6). One step of (stochastic) gradient descent can be used to update $\boldsymbol{\theta}$ based on (7).
III-B Mean-centered Feature Sparsification
@@
Update rule (7) shows $\overline{\boldsymbol{a}}$ approaches $\mathcal{H}_{2\lambda\gamma}(\overline{\boldsymbol{a}}_{k-1})$ by the second term while the weight for this term is increasing (equivalently $\lambda$ is decreasing) during the training. Based on this intuition, we propose to use the $\mathcal{H}_{2\lambda\gamma}(\overline{\boldsymbol{a}}_{k-1})$ in the forward propagation of the model. If we add the element-wise mean subtraction and addition in 2 into the hard-thresholding operator, we achieve the curve in the bottom of the first column in 1 assuming $\alpha\sigma_{\mathrm{ch}}=(2\lambda\gamma)^{2}$ (for simplicity of notation, we use $Th$ instead). This curve represents a sparsification around the feature mean value $\mu_{a}$. In other words, we have the following element-wise operation over the input $\boldsymbol{a}^{(\text{in})}$:
[…]
@@ III-C MeanSparse Design
III-C1 Sparsification in Post-processing
@@
The two approaches differ significantly: while training with MeanSparse can improve robustness by influencing learned representations, it requires a carefully designed threshold scheduler to avoid issues like gradient zeroing or instability, particularly in large models, with early experiments showing unstable training and failed convergence. In contrast, the post-training approach integrates more easily, leveraging established model statistics and requiring only a search over alpha values, making it scalable to large models. It has been successfully applied to architectures like Swin-L, achieving a $2.56\%$ robustness improvement without destabilization. Given the challenges of training MeanSparse in large models and the success of post-training integration, we chose the latter. This involves freezing the model, applying sparsification over mean-centered features before activation functions, calculating the mean value over the training set, and determining the optimal threshold ($Th$) to maximize robustness while maintaining clean accuracy.
III-C2 Adaptive Sparsification Using Feature Standard Deviation
@@
III-C3 Per-channel Sparsification
Throughout the neural network, the input to the activation functions can be considered as 4-dimensional, represented as $N\times C\times H\times W$ (NN, $C$, $H$ and $W$ represents the batch size, channels, height and width, respectively). To better capture the statistics in the representation, we use per-channel mean $\mu_{\mathrm{ch}}$ and variance $\sigma_{\mathrm{ch}}$ in sparsification operator. As a result, the mean and variance vectors are $C$-dimensional vectors. The sparsification operates on each feature channel separately, using the mean and variance of the corresponding channel.
III-C4 Complete Pipeline
The MeanSparse pipeline integrates mean-centered sparsification with three key modifications, placing it before all activation functions with a shared $\alpha$ parameter. Figure 1 shows an example applied to an adversarially trained RaWideResNet on the CIFAR-10 dataset. The first column illustrates a histogram of one channel’s features in the first layer, alongside its mean $\mu_{\mathrm{ch}}$ and standard deviation $\sigma_{\mathrm{ch}}$. The corresponding sparsification operator blocks variations near $\alpha\sigma_{\mathrm{ch}}$ around the mean. The second and third columns show the input and output of this operator for a test image from the CIFAR-10 dataset in sample space and histogram. Notably, the feature histogram is bimodal, with one mode aligning with $\mu_{\mathrm{ch}}$, representing uninformative variations blocked by the operator.